# arXiv:2602.23350v1[math.FA] 26 Feb 2026

### v1

@@ Abstract
We prove that if a sufficiently regular even log-concave measure satisfies a certain stronger form of the dimensional Brunn-Minkowski conjecture, then it also satisfies the (B)-conjecture. Furthermore, we show that hereditarily convex measures satisfy the aforementioned strengthened form, therefore providing an alternative proof of a recent result by Cordero-Erausquin and Eskenazis stating that a hereditarily convex measure satisfies both conjectures.
MSC 2020 Classification: 52A40, 52A20, 28C20, 46N20, 47F10.
@@ 1 Introduction
holds for every $t\in[0,1]$. Notice that (dim-BM) in the case of the Lebesgue measure reduces to the celebrated Brunn–Minkowski inequality (see, e.g., [13, Section 7]). The dimensional Brunn–Minkowski conjecture was first stated by Gardner and Zvavitch [8] for the Gaussian measure, a case settled by Eskenazis and Moschidis [6], while Cordero-Erausquin and Rotem [5] proved (dim-BM) for a broader class of measures, including sufficiently regular rotationally invariant log-concave measures. Nevertheless, the conjecture remains open for general log-concave measures. See [1, 4] for some recent developments and further references.
The (B)-conjecture, seemingly independent of dim-BM, concerns the behaviour of suitable dilates of convex bodies with respect to a log-concave measure. It asks whether, for an even log-concave measure $\mu$ on $\mathbb{R}^{n}$, the function
@@
is log-concave for every origin-symmetric convex $K\subset\mathbb{R}^{n}$. For the case of the Gaussian measure, this question was affirmatively answered by Cordero-Erausquin, Fradelizi, and Maurey [3]. Cordero-Erausquin and Rotem [5] later proved that the conjecture holds for the same general class of measures as (dim-BM) above. More recently, Cordero-Erausquin and Eskenazis [4] established the conjecture for so-called hereditarily convex measures (see later (4)), of which the measures treated in [5] are a subset.
In [4] it is shown that hereditarily convex measures satisfy both (dim-BM) and (B), extending earlier work of Kolesnikov and Livshyts [9], who proved that, for sufficiently regular log-concave measures, hereditary convexity implies (dim-BM). Although no direct implication between (dim-BM) and (B) is currently known, we show in this note that a suitable strengthening of (dim-BM) yields (B). Furthermore, we prove that this strengthening is itself implied by hereditary convexity, thereby providing an alternative route to some of the results in [4]. Our approach relies on a new method introduced in [1], combined with an application of the $L_{2}$ method developed by Kolesnikov and Milman [10, 11], which was in turn inspired by [3].
@@
Let $K\subset\mathbb{R}^{n}$ be a compact convex set with non-empty interior, such that its boundary $\partial K$ is a manifold of class $C^{2}$ with strictly positive principal curvatures. We shall say that such a set $K$ is of class $C^{2}_{+}$. Adopting the notation introduced by Livshyts [12], the concavity power of $K$ with respect to $\mu$ is the maximal exponent $p(\mu,K)$ such that
@@
where $K(\rho,t)$ denotes a perturbation of $K$ generated by a suitable function $\rho:\partial K\to\mathbb{R}$, and $\mathcal{A}$ is a class of admissible perturbations. We will elaborate more on this point in the next section. By a standard density argument, it can be shown that (dim-BM) holds for $\mu$ if and only if $p(\mu,K)\geq 1/n$ for every origin-symmetric $K$ of class $C^{2}_{+}$, see [12, Section 3.2] for a detailed exposition. Similarly, when studying the validity of (B), we may, by a standard density argument, restrict our attention to compact sets with sufficiently regular boundaries.
Given a non-empty compact convex set $K\subset\mathbb{R}^{n}$, we recall that its support function is defined as $h_{K}(x)=\sup_{y\in K}\langle x,y\rangle$. Moreover, if $K$ is of class $C^{2}_{+}$, we denote by $\nu_{K}$ its Gauss map. Throughout this note, given a compact convex set $K$ with non-empty interior and measure $\mu$ with continuous density with respect to the Lebesgue measure on $K$ (as will always be the case here), we ubiquitously use the symbol $\mathrm{d}\mu$ both for integration on $K$ and on $\partial K$. In the latter case, $\mu$ is understood as its restriction to the (n1)(n-1)-dimensional Hausdorff measure on $\partial K$. We are now ready to state our first main result.
@@
Let $u\in C^{2}(\mathbb{R}^{n})$ be an even convex function such that its Hessian matrix $\nabla^{2}u$ is positive definite, and let $\mu$ be the associated log-concave measure with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. Given a convex and compact set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, if
$\frac{\mu(K)}{\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu}\leq p(\mu,K),$ (3)
then
@@
In particular, if $\mu$ satisfies (3) for every $K$ of class $C^{2}_{+}$, $\mu$ satisfies the (B)-conjecture as well.
We say that the even log-concave measure $\mu$ satisfies the strong dimensional Brunn–Minkowski condition if (3) holds for every origin-symmetric $K$ of class $C^{2}_{+}$. The fact that this condition implies the dimensional Brunn–Minkowski conjecture (dim-BM) follows immediately from an integration by parts, together with the symmetry of the problem. See Corollary 2.2 later.
To state our second result, we introduce some additional notation. Given $u$ and $\mu$ as in Theorem 1.1, for every $\psi\in C^{2}(\mathbb{R}^{n})$ we define the $\mu$-weighted Laplacian as $\mathcal{L}_{\mu}(\psi)=\Delta\psi-\langle\nabla u,\nabla\psi\rangle$. Moreover, we say that a measure $\eta$ is log-concave with respect to $\mu$ if $\eta$ is absolutely continuous with respect to $\mu$ and its corresponding density $\mathrm{d}\eta/\,\mathrm{d}\mu$ is a log-concave function. The notion of hereditary convexity put forth in [4], inspired by [9] in the particular case of log-concave measures, is defined as follows. A log-concave measure $\mu$ is hereditarily convex if it is even, and for every even measure $\eta$ which is log-concave with respect to $\mu$ and every even $\psi$ such that $\mathcal{L}_{\mu}(\psi)\in L_{2}(\eta)$, one has
@@
Here, $||\cdot||_{\rm HS}$ denotes the Hilbert–Schmidt norm of matrices, and $|x|^{2}=\langle x,x\rangle$. Our second result reads as follows.
@@
To summarize this note, the results presented above show that if $\mu$ is an even-log concave measure as in Theorem 1.1, the implications in [4, (62)] can be refined as follows:
@@ 2 Preliminaries and proofs of the results
Before starting with the proofs, we introduce and recall some auxiliary notions and results. Consider a compact and convex set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. First, we elaborate on the concavity power defined in (2). There, the perturbations $K(\rho,t)$ are defined as follows: As $K$ is of class $C^{2}_{+}$, it is well-known that for every $\rho\in C^{2}(\partial K)$ there exists $\varepsilon>0$ such that $h_{t}=h_{K}+t\rho(\nu^{-1}_{K})$ are the support functions of compact and convex sets of class $C^{2}_{+}$ for every $|t|<\varepsilon$. We consider the sets $K(\rho,t)$ corresponding to the support functions $h_{t}$. Therefore, if $K$ is of class $C^{2}_{+}$, the family of admissible perturbations is, in fact, $C^{2}(\partial K)$. For more details on the theory of convex sets, the reader can consult the monograph by Schneider [13].
Given a compact and convex set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, we denote its second fundamental form by $\rm II$} and the tangential gradient on $\partial K$ by $\nabla_{\partial K}$. Consider, moreover, a convex function $u\in C^{2}(\mathbb{R}^{n})$ with strictly positive Hessian matrix and the measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. We write the weighted mean curvature of $\partial K$ as ${\rm H}_{\mu}=\mathrm{tr}({\rm II})-\langle\nabla u,\nu_{K}\rangle$. With a slight abuse of notation, we shall use $\langle\cdot,\cdot\rangle$ to denote also the restriction of the scalar product of $\mathbb{R}^{n}$ to the tangent spaces of $\partial K$. We write $H^{1}(\partial K,\mu)$ for the Sobolev space of functions on $\partial K$ with respect to the measure $\mu$. The following statement, proved in [1, Theorem 1.3], provides an analytic representation for the concavity power $p(\mu,K)$.
@@
Consider $\mu$ such that it satisfies the assumptions of Theorem 1.1, and a compact and convex set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. Then there exists a unique weak solution $\overline{\rho}\in H^{1}(\partial K,\mu)$ to the equation
@@
Going back to condition (3), observe that integration by parts yields the identity
@@
Consider a log-concave measure $\mu$ as in Theorem 1.1. If $\mu$ satisfies (3) for every compact and convex set of class $C^{2}_{+}$, then (dim-BM) holds true.
@@
Then, $\mathcal{E}$ is a uniformly elliptic operator on $\partial K$ and $\overline{\rho}$ is a weak solution of $\mathcal{E}(\rho)=1$. We remark that in the case that the support function of $K$ is assumed to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, all the coefficients in (9) are of class at least $C^{1}$. Under this assumption, the function $\overline{\rho}$ is, in fact, a classical solution for (5), as can be seen through classical elliptic regularity results. See, e.g., [7]. Now, observe that integration by parts yields the identity
@@
The right-hand side is positive thanks to the weighted Poincaré inequality (compare [11]), which can be obtained by direct differentiation of (1). Therefore,
@@
for every $\rho\in C^{2}(\partial K)$. We remark that this inequality holds for less regular functions. In fact, it remains true for functions in $H^{1}(\partial K,\mu)$, as proved in [1, Lemma 3.1]. However, in this note, we will always assume that the functions involved are at least $C^{2}$. Additionally, we remark that the action of $\mathcal{E}$ on the support function, whenever $h_{K}$ is of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, is
@@
Next, we recall a local formulation of the (B)-conjecture. Differentiating twice $\mu(e^{t}K)$ for $\mu$ and $K$ as above (see, for example, [5, (10)], which in turn generalizes [3, (4)]), it follows that (B) is equivalent to
$\frac{\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu}{\mu(K)}-\left(\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}\right)^{2}\leq\frac{\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu}{\mu(K)}+\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}$ (12)
for every compact and origin-symmetric convex set $K\subset\mathbb{R}^{n}$ with non-empty interior. For the $\mu$-weighted Laplacian $\mathcal{L}_{\mu}$, we recall the following useful integration by parts identity, which holds for any two functions $\varphi,\psi\in C^{2}(K)$,
@@
Finally, we will make use of the following Reilly formula, proved in [10, (2.3)] in the more general context of weighted Riemannian manifolds (compare also [4, (21)]). Consider a function $\psi\in C^{2}(\mathbb{R}^{n})$ and its normal derivative $\rho=\langle\nabla\psi,\nu_{K}\rangle$. Then
@@
We begin by following the argument in [1, Proposition 5.3]. Let $\overline{\rho}$ be given by (5). We may assume by approximation that, in addition, the support function of $K$ is of class $C^{3}$ and thus $\overline{\rho}\in C^{2}(\partial K)$. Moreover, denote by $\eta$ the measure with density $\mathrm{d}\eta=\frac{1}{\mu(K)}\mathbbm{1}_{K}\,\mathrm{d}\mu$, where $\mathbbm{1}_{K}$ is the characteristic function of $K$ such that $\mathbbm{1}_{K}(x)=1$ if $x\in K$ and 0 otherwise. From (11) and since $\mathcal{E}(\overline{\rho})=1$ we infer
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\int_{\partial K}(h_{K}-\overline{\rho})\mathcal{E}(\overline{\rho})\,\mathrm{d}\eta=\int_{\partial K}\mathcal{E}(h_{K}-\overline{\rho})\overline{\rho}\,\mathrm{d}\eta$
@@
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\underbrace{\int_{\partial K}f\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}f\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm A})}$
$+\underbrace{\int_{\partial K}h_{K}(\nu_{K})\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm B})}$ (15)
We begin by computing the term $({\rm A})$. Notice that by (11) and (5) we have
@@
$({\rm A})=-\int_{\partial K}f\mathcal{E}(f)\,\mathrm{d}\eta\leq 0\,$,
where the last inequality is (10). We proceed with the term $({\rm B})$. Notice that since for every $x\in\partial K$ we have that $h_{K}(\nu_{K}(x))=\langle x,\nu_{K}(x)\rangle=\langle\nabla(|x|^{2}/2),\nu_{K}(x)\rangle$, an application of (13) implies
@@
Moreover, integrating by parts again shows
@@
$({\rm B})=\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\eta+\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta\right)^{2}.$ (17)
If the strong dimensional Brunn–Minkowski holds, by (7) we have that $({\rm A})+({\rm B})\geq 0$ and, since $({\rm A})\leq 0$, we conclude that $({\rm B})\geq 0$. That is,
@@
which is precisely (12). Thus, (B) holds, concluding the proof.
@@
For this proof, we apply Hörmander’s $L_{2}$ method, following the approach introduced in [10, 11], which can be consulted for further details. See also the discussion in [4, Section 2]. Consider the function $\overline{\rho}$ given by (5). Let $\psi$ be the unique solution to the Neumann problem
@@
Such solution exists provided that the integral (with respect to $\mu$) of the interior data on $K$ is the same as the integral of the boundary data on $\partial K$, which is indeed the case here. By approximation, we can assume $h_{K}$ to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, so that $\overline{\rho}\in C^{2}(\partial K)$. Compare the discussion after (9). Thus, standard results in elliptic regularity theory (see, for example, [7, Section 6]) yield that $\psi\in C^{2}(K)$. Consider the normalized measure $\eta$ as in the previous proof. The Cauchy–Schwarz inequality for positive quadratic forms implies that
@@
which, in combination with (19),(20) and the fact that $\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta>0$ implies
@@
Thus, the strong dimensional Brunn–Minkowski conjecture holds for the measure $\mu$, concluding the proof.

### v2

# A strengthening of the dimensional Brunn–Minkowski conjecture implies the (B)-conjecture

@@ Abstract
In this note, we prove that if a sufficiently regular even log-concave measure satisfies a certain stronger form of the dimensional Brunn-Minkowski conjecture, then it also satisfies the (B)-conjecture. Furthermore, we show that hereditarily convex measures satisfy the aforementioned strengthened form, therefore providing an alternative proof of a recent result by Cordero-Erausquin and Eskenazis stating that a hereditarily convex measure satisfies both conjectures.
MSC 2020 Classification: 52A40, 52A20, 46N20, 47F10.
@@ 1 Introduction
holds for every $t\in[0,1]$. Notice that (dim-BM) in the case of the Lebesgue measure reduces to the celebrated Brunn–Minkowski inequality (see, e.g., [16, Section 7]), which does not require any symmetry for $K$ and $L$. The dimensional Brunn–Minkowski conjecture was first stated by Gardner and Zvavitch [9] for the Gaussian measure, a case settled by Eskenazis and Moschidis [7], while Cordero-Erausquin and Rotem [6] proved (dim-BM) for a broader class of measures, including sufficiently regular rotationally invariant log-concave measures. Nevertheless, the conjecture remains open for general even log-concave measures. See [1, 5] for some recent developments and further references.
The (B)-conjecture, seemingly independent of (dim-BM), concerns the behaviour of suitable dilates of convex bodies with respect to a log-concave measure. It asks whether, for an even log-concave measure $\mu$ on $\mathbb{R}^{n}$, the function
@@
is log-concave for every origin-symmetric convex $K\subset\mathbb{R}^{n}$. This problem was put forth by Banaszczyk and popularised by Latała [14]. For the case of the Gaussian measure, this question was affirmatively answered by Cordero-Erausquin, Fradelizi, and Maurey [4]. A previous proof by Brascamp and Lieb [3], which remained unnoticed for decades, recently resurfaced in [5]. Cordero-Erausquin and Rotem [6] later proved that the conjecture holds for the same general class of measures as (dim-BM) above. More recently, Cordero-Erausquin and Eskenazis [5] established the conjecture for so-called hereditarily convex measures (see later (4)), of which the measures treated in [6] are a subset.
In [5] it is shown that hereditarily convex measures satisfy both (dim-BM) and (B), extending earlier work of Kolesnikov and Livshyts [10], who proved that, for sufficiently regular log-concave measures, hereditary convexity implies (dim-BM). Although no direct implication between (dim-BM) and (B) is currently known, we show in this note that a suitable strengthening of (dim-BM) yields (B). Furthermore, we prove that this strengthening is implied by hereditary convexity, thereby providing an alternative route to some of the results in [5]. Our approach relies on a new method introduced in [1], combined with an application of the $L_{2}$ method developed by Kolesnikov and Milman [11, 12], which was in turn inspired by [4].
@@
Let $K\subset\mathbb{R}^{n}$ be a compact and convex set with non-empty interior, such that its boundary $\partial K$ is a manifold of class $C^{2}$ with strictly positive principal curvatures. We shall say that such a set $K$ is of class $C^{2}_{+}$. Adopting the notation introduced by Livshyts [15], the concavity power of $K$ with respect to $\mu$ is the maximal exponent $p(\mu,K)$ such that
@@
where $K(\rho,t)$ denotes a perturbation of $K$ generated by a suitable function $\rho:\partial K\to\mathbb{R}$, and $\mathcal{A}$ is a class of admissible perturbations. We will elaborate more on this point in the next section. By a standard density argument, it can be shown that (dim-BM) holds for $\mu$ if and only if $p(\mu,K)\geq 1/n$ for every origin-symmetric $K$ of class $C^{2}_{+}$, see [15, Section 3.2] for a detailed exposition. Similarly, when studying the validity of (B), we may, by a standard density argument, restrict our attention to compact sets with sufficiently regular boundaries.
Given a non-empty compact and convex set $K\subset\mathbb{R}^{n}$, we recall that its support function is defined as $h_{K}(x)=\sup_{y\in K}\langle x,y\rangle$. Moreover, if $K$ is of class $C^{2}_{+}$, we denote by $\nu_{K}:\partial K\to\mathbb{S}^{n-1}$ its Gauss map. Throughout this note, given a compact and convex set $K$ with non-empty interior and a measure $\mu$ with continuous density with respect to the Lebesgue measure on $K$ (as will always be the case here), we ubiquitously use the symbol $\mathrm{d}\mu$ both for integration on $K$ and on $\partial K$. In the latter case, $\mu$ is understood as its restriction to the (n1)(n-1)-dimensional Hausdorff measure on $\partial K$. We are now ready to state our first main result.
@@
Let $u\in C^{2}(\mathbb{R}^{n})$ be an even function such that $\nabla^{2}u$ is positive definite, and let $\mu$ be the associated log-concave measure with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. Suppose that for every origin-symmetric set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$,
$\frac{\mu(K)}{\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu}\leq p(\mu,K).$ (3)
Then, for every compact and convex and origin-symmetric set $K$
@@
In other words, if $\mu$ satisfies (3) for every origin-symmetric $K$ of class $C^{2}_{+}$, $\mu$ satisfies the (B)-conjecture as well.
We say that the even log-concave measure $\mu$ satisfies the strong dimensional Brunn–Minkowski condition if (3) holds for every origin-symmetric set $K$ of class $C^{2}_{+}$. The fact that this condition implies the dimensional Brunn–Minkowski conjecture (dim-BM) follows immediately from an integration by parts, together with the symmetry of the problem. See Corollary 2.2 later.
To state our second result, we introduce some additional notation. Given $u$ and $\mu$ as in Theorem 1.1, for every $\psi\in C^{2}(\mathbb{R}^{n})$ we define the $\mu$-weighted Laplacian as $\mathcal{L}_{\mu}(\psi)=\Delta\psi-\langle\nabla u,\nabla\psi\rangle$. Moreover, we say that a measure $\eta$ is log-concave with respect to $\mu$ if $\eta$ is absolutely continuous with respect to $\mu$ and its corresponding density $\mathrm{d}\eta/\,\mathrm{d}\mu$ is a log-concave function. The notion of hereditary convexity was formally introduced in [5]. We restrict our treatment to the case of log-concave measures: A log-concave measure $\mu$ is hereditarily convex if it is even, and for every even measure $\eta$ which is log-concave with respect to $\mu$ and every even $\psi$ such that $\mathcal{L}_{\mu}(\psi)\in L_{2}(\eta)$, one has
@@
Here, $||\cdot||_{\rm HS}$ denotes the Hilbert–Schmidt norm of matrices, and $|x|^{2}=\langle x,x\rangle$. As previously mentioned, it is known that the class of hereditarily convex measures includes radially symmetric log-concave measures. It remains an open question of major interest whether every even log-concave measure is hereditarily convex; compare [5, Question 16]. Our second result reads as follows.
@@
To summarize this note, the results presented above show that if $\mu$ is an even-log concave measure as in Theorem 1.1, the implications in [5, (62)] can be refined as follows:
@@ 2 Preliminaries and proofs of the results
Before starting with the proofs, we introduce and recall some auxiliary notions and results. Consider a compact and convex set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. First, we elaborate on the concavity power defined in (2). There, the perturbations $K(\rho,t)$ are defined as follows: As $K$ is of class $C^{2}_{+}$, it is well-known that for every $\rho\in C^{2}(\partial K)$ there exists $\varepsilon>0$ such that $h_{t}=h_{K}+t\rho(\nu^{-1}_{K})$ are the support functions of compact and convex sets of class $C^{2}_{+}$ for every $|t|<\varepsilon$. We consider the sets $K(\rho,t)$ corresponding to the support functions $h_{t}$. Therefore, if $K$ is of class $C^{2}_{+}$, the family of admissible perturbations is, in fact, $C^{2}(\partial K)$. See also [13, Section 3.2] and, for more details on the theory of convex sets, the monograph by Schneider [16].
Given a set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, we denote its second fundamental form by $\rm II$} and the tangential gradient on $\partial K$ by $\nabla_{\partial K}$. Consider, moreover, a function $u\in C^{2}(\mathbb{R}^{n})$ with strictly positive Hessian matrix and the measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. We write the weighted mean curvature of $\partial K$ as ${\rm H}_{\mu}=\mathrm{tr}({\rm II})-\langle\nabla u,\nu_{K}\rangle$. With a slight abuse of notation, we shall use $\langle\cdot,\cdot\rangle$ to denote also the restriction of the scalar product of $\mathbb{R}^{n}$ to the tangent spaces of $\partial K$. We write $H^{1}(\partial K,\mu)$ for the Sobolev space of functions on $\partial K$ with respect to the measure $\mu$. The following statement, proved in [1, Theorem 1.3], provides an analytic representation for the concavity power $p(\mu,K)$.
@@
Consider $\mu$ such that it satisfies the assumptions of Theorem 1.1, and a compact and convex origin-symmetric set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. Then there exists a unique weak solution $\overline{\rho}\in H^{1}(\partial K,\mu)$ to the equation
@@
Going back to condition (3), observe that integration by parts yields the identity (see later (16) for a proof)
@@
Consider a log-concave measure $\mu$ as in Theorem 1.1. If $\mu$ satisfies (3) for every compact, convex and origin-symmetric set of class $C^{2}_{+}$, then (dim-BM) holds true.
@@
Then, $\mathcal{E}$ is a uniformly elliptic operator on $\partial K$ and $\overline{\rho}$ is a weak solution of $\mathcal{E}(\rho)=1$. We remark that if the support function of $K$ is assumed to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, all the coefficients in (9) are of class at least $C^{1}$. Under this assumption, the function $\overline{\rho}$ is, in fact, a classical solution for (5), as can be seen through classical elliptic regularity results. See, e.g., [8]. Now, observe that integration by parts yields the identity
@@
The right-hand side is non-negative thanks to the weighted Poincaré inequality (compare [12]), which can be obtained by direct differentiation of (1). Therefore,
@@
for every $\rho\in C^{2}(\partial K)$. We remark that this inequality holds for less regular functions. Additionally, we remark that the action of $\mathcal{E}$ on the support function, whenever $h_{K}$ is of class $C^{2}$ in $\mathbb{R}^{n}\setminus\{0\}$, is
@@
Next, we recall a local formulation of the (B)-conjecture. Differentiating twice $\mu(e^{t}K)$ for $\mu$ and $K$ as above (see, for example, [6, (10)], which in turn generalizes [4, (4)]), shows that the log-concavity of the function $t\mapsto\mu(e^{t}K)$ at $t=0$ is equivalent to
$\frac{\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu}{\mu(K)}-\left(\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}\right)^{2}\leq\frac{\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu}{\mu(K)}+\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}.$ (12)
Thus, (B) is equivalent to the validity of (12) for every compact and origin-symmetric convex set $K\subset\mathbb{R}^{n}$ with non-empty interior. For the $\mu$-weighted Laplacian $\mathcal{L}_{\mu}$, we recall the following useful integration by parts identity, which holds for any two functions $\varphi,\psi\in C^{2}(K)$:
@@
Finally, we will make use of the following Reilly formula, proved in [11, (2.3)] in the more general context of weighted Riemannian manifolds (compare also [5, (21)]). Consider a function $\psi\in C^{2}(\mathbb{R}^{n})$ and its normal derivative $\rho=\langle\nabla\psi,\nu_{K}\rangle$. Then
@@
We begin by following the argument in [1, Proposition 5.3]. Let $K\subset\mathbb{R}^{n}$ be a compact and convex origin-symmetric set and let $\overline{\rho}$ be given by (5). We may assume by approximation that, in addition, the support function of $K$ is of class $C^{3}$ with strictly positive principal curvatures and, thus, $\overline{\rho}\in C^{2}(\partial K)$. Moreover, denote by $\eta$ the measure with density $\mathrm{d}\eta=\frac{1}{\mu(K)}\mathbbm{1}_{K}\,\mathrm{d}\mu$, where $\mathbbm{1}_{K}$ is the characteristic function of $K$ such that $\mathbbm{1}_{K}(x)=1$ if $x\in K$ and 0 otherwise. Suppose that (3) holds, we shall show that (12) is also satisfied. From (11) and since $\mathcal{E}(\overline{\rho})=1$ we infer
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\int_{\partial K}(h_{K}(\nu_{K})-\overline{\rho})\mathcal{E}(\overline{\rho})\,\mathrm{d}\eta=\int_{\partial K}\mathcal{E}(h_{K}(\nu_{K})-\overline{\rho})\overline{\rho}\,\mathrm{d}\eta$
@@
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\underbrace{\int_{\partial K}f\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}f\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm I})}$
$+\underbrace{\int_{\partial K}h_{K}(\nu_{K})\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm I\!I})}.$ (15)
We begin by computing the term $({\rm I})$. Notice that by (11) and (5) we have
@@
$({\rm I})=-\int_{\partial K}f\mathcal{E}(f)\,\mathrm{d}\eta\leq 0\,$,
where the last inequality is (10). We proceed with the term $({\rm I\!I})$. Notice that since for every $x\in\partial K$ we have that $h_{K}(\nu_{K}(x))=\langle x,\nu_{K}(x)\rangle=\langle\nabla(|x|^{2}/2),\nu_{K}(x)\rangle$, an application of (13) implies
@@
Moreover, integrating again by parts shows
@@
$({\rm I\!I})=\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\eta+\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta\right)^{2}.$ (17)
If the strong dimensional Brunn–Minkowski condition holds, by (7) we have that $({\rm I})+({\rm I\!I})\geq 0$ and, since $({\rm I})\leq 0$, we conclude that $({\rm I\!I})\geq 0$. That is,
@@
which is precisely (12). Since $K$ was arbitrary we infer that (B) holds, concluding the proof.
@@
For this proof, we apply Hörmander’s $L_{2}$ method, following the approach introduced in [11, 12], which can be consulted for further details. See also the discussion in [5, Section 2]. Given $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, we shall prove that (3) is satisfied. Consider the function $\overline{\rho}$ solving (5) for the previous choice of $K$, and let $\psi$ be the unique solution to the Neumann problem
@@
Such solution exists provided that the integral (with respect to $\mu$) of the interior data on $K$ is the same as the integral of the boundary data on $\partial K$, which is indeed the case here. By approximation, we can assume $h_{K}$ to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, so that $\overline{\rho}\in C^{2}(\partial K)$. Compare the discussion after (9). Thus, standard results in elliptic regularity theory (see, for example, [11, Theorem 2.5]) yield that $\psi\in C^{2}(K)$. Consider the normalized measure $\eta$ as in the previous proof. The Cauchy–Schwarz inequality for positive quadratic forms implies that
@@
which, in combination with (19), (20), and the fact that $\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta>0$ implies
@@
Thus, (3) holds for an arbitrary compact, convex and origin-symmetric set $K$ and the strong dimensional Brunn–Minkowski condition holds for the measure $\mu$, concluding the proof.